OpenStudy (anonymous):

Find two consecutive even integers such that three times the lesser integer, added to the larger integer, is 58.

Parth (parthkohli):

3x + x + 1 = 58 => 4x + 1 = 58 => 4x = 57 => x = 19 So, the two integers are 19 and 19+1 = 20

OpenStudy (anonymous):

So is the final answer 20?

OpenStudy (anonymous):

Im confused?

Parth (parthkohli):

The two integers are 19 and 20

OpenStudy (lgbasallote):

@ParthKohli seems to have overlooked consecutive EVEN numbers

Parth (parthkohli):

OOPS. yes, thanks for correcting

OpenStudy (anonymous):

So 20 isn't the answer then?

Parth (parthkohli):

No

OpenStudy (lgbasallote):

my work here is done <tips imaginary hat>

Parth (parthkohli):

3x + x + 2 = 58 4x = 56 x = 14 Answer: integers are 14 and 16

Parth (parthkohli):

Medal please?